### v1

# Kernels of paired operators and their adjoints

Abstract
We review the basic properties of paired operators and their adjoints, the transposed paired operators, with particular reference to commutation relations, and we study the properties of their kernels, bringing out their similarities and also, somewhat surprisingly, their stark differences. Various notions expressing different invariance properties are also reviewed and we extend to paired operators some known invariance results.
Keywords: paired operator; transposed paired operator; Toeplitz operator; invariant subspace; almost-invariant subspace; nearly-invariant subspace; truncated Toeplitz operator; kernel.
MSC (2010): 30H10, 47B35, 47B38
This work was partially supported by FCT/Portugal through CAMGSD, IST-ID, projects UIDB/04459/2020 and UIDP/04459/2020.
1 Introduction
The canonical example of a singular integral equation on $L^{p}(\mathbb{T})$, $1<p<\infty$, is
$a(t)f(t)+b(t)(S_{\mathbb{T}}f)(t)=g(t)$ (1.1)
where $a$ and $b$ are bounded functions in $\mathbb{T}$, $g$ is a given function in $L^{p}(\mathbb{T})$ and $S_{\mathbb{T}}$ is the singular integral operator with Cauchy kernel defined by
[…]
The boundedness, Fredholmness and invertibility of the operator
$T=aI+bS_{\mathbb{T}}$ (1.3)
defined by the left hand side of (1.1), as well as methods to obtain explicit solutions to that equation and systems of equations of the same type, have been extensively studied, mainly in connection with factorisation of Wiener–Hopf type (see for example [30, 14, 21] and references therein). In general certain conditions are imposed on $a$ and $b$, such as $a^{2}-b^{2}$ being bounded away from zero [30, 21]. For $p=2$, it was Shinbrot who first derived a constructive general method to find explicit solutions to (1.1), by placing it in an abstract Hilbert space setting, in [28], using the fact that
$I=P^{+}+P^{-}\qquad\hbox{and}\qquad S_{\mathbb{T}}=P^{+}-P^{-},$ (1.4)
where $P^{\pm}$ are the orthogonal projections from $L^{2}$ onto $H_{+}^{2}:=H^{2}(\mathbb{D})$ and $H_{-}^{2}=(H_{+}^{2})^{\perp}$, respectively. Indeed (1.4) implies that (1.3) can be written as a paired operator [21]
$S_{A,B}=AP+BQ$ (1.5)
where $A$ and $B$ are bounded operators and $P$, $Q$ are complementary projections on a Banach space. In [28] Shinbrot assumed that $A^{-1}$ existed and $A^{-1}B$ was selfadjoint and of single sign, noting that the situation would be much more complicated in the general case.
Naturally, one can also consider singular integral equations of the form
$af+bS_{\mathbb{T}}f=g$ (1.6)
to which, by (1.4), one can associate in a similar way another paired operator of the form
$\Sigma_{A,B}=P^{+}A+P^{-}B,$ (1.7)
which we call a transposed paired operator.
If $A$ and $B$ are invertible operators, then (1.5) and (1.7) are said to be of normal type [21]. This is by far the most studied case. Then we have that any operator of the form (1.7) is equivalent to an operator of the form (1.5), i.e., they are related by products with invertible operators:
$PA+QB=[(I+PAB^{-1}Q)A](B^{-1}P+A^{-1}Q)[(I-QAB^{-1}P)B].$ (1.8)
Moreover, if $AB=BA$, then [21, III, Theorem 1.2]
$PA+QB=[(I+PAB^{-1}Q)B^{-1}](AP+BQ)[(I-QAB^{-1}P)B].$ (1.9)
This is true, in particular, if $A$ and $B$ are multiplication operators on $L^{2}$, as it happens in the case of singular integral operators.
Thus, for many purposes (such as the study of boundedness, Fredholmness, invertibility, solvability of equations) it suffices to consider only one type of paired operators. This may explain why transposed paired operators have hardly been considered separately in the literature on this topic, in spite of renewed interest in recent years [15, 29, 22, 23, 11, 4, 8].
In this paper we shall work in the context $X=L^{2}(\mathbb{T})$, where $\mathbb{T}$ is the unit circle, and $P=P^{+}$, $Q=P^{-}$. For functions $a,b\in L^{\infty}(\mathbb{T})=:L^{\infty}$, we take $A=M_{a}$ and $B=M_{b}$, the corresponding multiplication operators. Then we write $S_{a,b}$ and $\Sigma_{a,b}$ for the corresponding paired and transposed paired operators, i.e.,
$S_{a,b}=aP^{+}+bP^{-}\qquad\hbox{and}\qquad\Sigma_{a,b}=P^{+}aI+P^{-}bI.$ (1.10)
Writing $T_{a}$ for the Toeplitz operator $P^{+}aP^{+}$ on $H^{2}_{+}$, we see that
$P^{+}S_{a,b}{}_{|H^{2}_{+}}=P^{+}\Sigma_{a,b}{}_{|H^{2}_{+}}=T_{a},$
so that paired operators are dilations of Toeplitz operators. Moreover, if $b\in GL^{\infty}$ (that is, invertible in $L^{\infty}$), then $S_{a,b}$ and $\Sigma_{a,b}$ are equivalent, and both are equivalent after extension to $T_{a/b}$ [3] (see also [1] for more on these concepts). Hence we have an isomorphism of kernels:
$\hbox{ker}\Sigma_{a,b}\approx\hbox{ker}S_{a,b}\approx\hbox{ker}T_{a/b}.$ (1.11)
The papers [4] and [8] were mainly concerned with paired operators $S_{a,b}$, but the theory of transposed paired operators is also very rich, and we shall also be discussing these, bringing out some of the similarities and differences.
We shall assume throughout the paper that $a,b\in L^{\infty}$ and
$a\neq 0,\quad b\neq 0\quad\hbox{a.e. on }\mathbb{T},\quad\hbox{and}\quad a=b\quad\hbox{or}\quad a-b\neq 0\quad\hbox{a.e. on }\mathbb{T}.$ (1.12)
If $a\neq b$ then we say that the pair $(a,b)$ is nondegenerate.
After describing some of the basic properties of paired operators in Section 2 we shall present some of the properties of paired kernels in Section 3, before moving on to transposed paired kernels in Section 4, in which the similarities and differences between the two notions will be identified. Finally, Section 5 is devoted to reviewing various notions expressing different invariance properties, such as almost-invariant and nearly-invariant subspaces for a given operator, analysing their relations and extending some known invariance results to kernels of paired operators. Along the way we get new insights into Toeplitz and Hankel operators and new formulations of known results for Toeplitz kernels.
Some notation: we shall write $\tilde{T}_{a}$ for the dual Toeplitz operator $P^{-}aP^{-}$ on $H_{-}^{2}$ and $\tilde{H}_{a}$ for the Hankel operator $P^{+}aP^{-}:H_{-}^{2}\rightarrow H_{+}^{2}$. In addition, we shall say that $O_{-}\in H_{-}^{2}$ (resp. $\overline{H^{\infty}}$) is outer if $\overline{z}\overline{O_{-}}$ (resp. $\overline{O}_{-}$) is outer in $H_{+}^{2}$ (resp. $H^{\infty}$).
2 Basic properties
We start by noting that the two types of paired operator, $S_{a,b}$ and $\Sigma_{a,b}$, are closely connected. Indeed, if $b\in\mathcal{G}L^{\infty}$, then $S_{a,b}$ and $\Sigma_{a,b}$ are equivalent, i.e., they can be related by invertible operators
$\Sigma_{a,b}=ES_{a,b}F$ (2.1)
with $E,F$ invertible [21, Ch. III, Sec. 1].
Moreover, for any $a,b\in L^{\infty}$ we have
$S_{a,b}^{*}=\Sigma_{\overline{a},\overline{b}}$ (2.2)
and
$\overline{S_{a,b}f}=zS_{\overline{b},\overline{a}}\overline{z}\overline{f}=z\Sigma^{*}_{b,a}(\overline{z}\overline{f}),$ (2.3)
so that some properties of $\Sigma_{a,b}$ can be deduced easily from the corresponding properties of $S_{a,b}$. For example, we have the following optimal estimates from [4].
Theorem 2.1.
For $a,b\in L^{\infty}$, let $m=\max(\|a\|_{\infty},\|b\|_{\infty})$ and $M=\|a\|_{\infty}+\|b\|_{\infty}$. Then
$m\leq\|S_{a,b}\|=\|\Sigma_{\overline{a},\overline{b}}\|\leq\min(M,\sqrt{2}\,m).$
Clearly, the same inequalities hold for $\Sigma_{a,b}$. As an immediate consequence we have:
Corollary 2.2.
$S_{a,b}=0\iff\Sigma_{a,b}=0\iff a=b=0,$ (2.4)
and it follows that the symbol pair $(a,b)$ is unique for any operator of the form (1.10).
Since, for any $\eta\in L^{\infty}$,
$M_{\eta}=S_{\eta,\eta}=\Sigma_{\eta,\eta},$
we have that $S_{a,b}$ (or $\Sigma_{a,b}$) is a multiplication operator on $L^{2}$ if and only if $a=b$.
In the same way that nonzero Toeplitz operators cannot be compact (let alone finite rank), a similar result holds for paired operators, and hence for transposed paired operators.
Proposition 2.3.
The operators $S_{a,b}$ and $\Sigma_{a,b}$ can only be compact if $a=b=0$.
Proof.
Note that $\Sigma_{a,b}=P^{+}aI+P^{-}bI$, and this can only be compact if the operators $P^{+}aI$ and $P^{-}bI$ are both compact, because a Hilbert space sequence $(x_{n}+y_{n})$ with $\langle x_{n},y_{m}\rangle=0$ for all $n$ and $m$ converges if and only if the two sequence $(x_{n})$ and $(y_{n})$ both converge.
Now $P^{+}aI:L^{2}\to H^{2}_{+}$ is an extension of the Toeplitz operator $T_{a}$ on $H^{2}_{+}$ and so is not compact unless $a=0$. Similarly for $P^{-}bI$. Hence $\Sigma_{a,b}$ is not compact unless $a=b=0$, and so neither is $S_{a,b}=\Sigma_{\overline{a},\overline{b}}^{*}$. ∎
We shall require certain commutation results from [4, Prop. 3.1]. Since for two operators on a Hilbert space we have $XT=TX$ if and only if $T^{*}X^{*}=X^{*}T^{*}$, and indeed the adjoint of the commutator $[X,T]=XT-TX$ is the commutator $[T^{*},X^{*}]$, which has the same rank, we can immediately apply the results we quote to the context of transposed paired kernels. Similarly, questions on the commutant of two Toeplitz operators, as analysed by Brown and Halmos in [2], have their generalizations to paired operators. The following can be found in [4].
Proposition 2.4.
For any nondegenerate $(a,b)$, $(\tilde{a},\tilde{b})$ satisfying (1.12), we have:
(i) $S_{a,b}S_{\tilde{a},\tilde{b}}=S_{c,d}\iff\tilde{a}\in H^{\infty}\hbox{ and }\tilde{b}\in\overline{H^{\infty}}$, and in that case $c=a\tilde{a}$ and $d=b\tilde{b}$;
(ii) $\Sigma_{a,b}\Sigma_{\tilde{a},\tilde{b}}=\Sigma_{c,d}\iff a\in\overline{H^{\infty}}\hbox{ and }b\in H^{\infty}$, and in that case $c=a\tilde{a}$ and $d=b\tilde{b}$.
As an immediate consequence of Proposition 2.4 we see the following.
Corollary 2.5.
For nondegenerate $(a,b)$ and $(\tilde{a},\tilde{b})$ we have $S_{a,b}S_{\tilde{a},\tilde{b}}=0$ if and only if $\tilde{a}\in H^{\infty}$, $\tilde{b}\in\overline{H^{\infty}}$ and $a\tilde{a}=b\tilde{b}=0$.
More generally, we have:
Proposition 2.6.
(i) $S_{a,b}S_{\tilde{a},\tilde{b}}-S_{a\tilde{a},b\tilde{b}}=(a-b)(P^{+}\tilde{b}P^{-}-P^{-}\tilde{a}P^{+})$;
(ii) $\Sigma_{a,b}\Sigma_{\tilde{a},\tilde{b}}-\Sigma_{a\tilde{a},b\tilde{b}}=(P^{-}bP^{+}-P^{+}aP^{-})(\tilde{a}-\tilde{b})I$.
Hence, regarding the commutation relations for two paired operators or transposed paired operators, we have the following.
Proposition 2.7.
(i) $S_{a,b}S_{\tilde{a},\tilde{b}}-S_{\tilde{a},\tilde{b}}S_{a,b}=\\ \-\qquad\qquad(a-b)(P^{+}\tilde{b}P^{-}-P^{-}\tilde{a}P^{+})-(\tilde{a}-\tilde{b})(P^{+}bP^{-}-P^{-}aP^{+})$;
(ii) $\Sigma_{a,b}\Sigma_{\tilde{a},\tilde{b}}-\Sigma_{\tilde{a},\tilde{b}}\Sigma_{a,b}=\\ \-\qquad\qquad(P^{+}\tilde{a}P^{-}-P^{-}\tilde{b}P^{+})(a-b)I-(P^{+}aP^{-}-P^{-}bP^{+})(\tilde{a}-\tilde{b})I$.
The right hand sides of the equations in Propositions 2.6 and 2.7 are expressed in terms of operators of Hankel type, which allows us to derive from these the following compactness results, where we write $\mathcal{C}$ for $C(\mathbb{T})$ and $\mathcal{R}$ for the space of all rational functions in $L^{\infty}(\mathbb{T})$.
Proposition 2.8.
Suppose that $\tilde{a}\neq\tilde{b}$. Then:
(i) $S_{a,b}S_{\tilde{a},\tilde{b}}-S_{a\tilde{a},b\tilde{b}}$ is a compact (resp. finite rank) operator if $\tilde{a}\in H^{\infty}+\mathcal{C}$, $\tilde{b}\in\overline{H^{\infty}}+\mathcal{C}$ (resp. $\tilde{a}\in H^{\infty}+\mathcal{R}$, $\tilde{b}\in\overline{H^{\infty}}+\mathcal{R}$).
(ii) $S_{a,b}S_{\tilde{a},\tilde{b}}-S_{\tilde{a},\tilde{b}}S_{a,b}$ is a compact (resp. finite rank operator) if $a,\tilde{a}\in H^{\infty}+\mathcal{C}$ and $b,\tilde{b}\in\overline{H^{\infty}}+\mathcal{C}$ (resp. $a,\tilde{a}\in H^{\infty}+\mathcal{R}$ and $b,\tilde{b}\in\overline{H^{\infty}}+\mathcal{R}$).
(iii) $S_{a,b}\eta I-\eta S_{a,b}$ is compact (resp. finite rank) if $\eta\in\mathcal{C}$ (resp. $\eta\in\mathcal{R}$).
The analogous results for transposed paired operators are the following.
Proposition 2.9.
Suppose that $\tilde{a}\neq\tilde{b}$. Then:
(i) $\Sigma_{a,b}\Sigma_{\tilde{a},\tilde{b}}-\Sigma_{a\tilde{a},b\tilde{b}}$ is a compact (resp. finite rank) operator if $\tilde{a}\in\overline{H^{\infty}}+\mathcal{C}$, $\tilde{b}\in H^{\infty}+\mathcal{C}$ (resp. $\tilde{a}\in\overline{H^{\infty}}+\mathcal{R}$, $\tilde{b}\in H^{\infty}+\mathcal{R}$).
(ii) $\Sigma_{a,b}\Sigma_{\tilde{a},\tilde{b}}-\Sigma_{\tilde{a},\tilde{b}}\Sigma_{a,b}$ is a compact (resp. finite rank operator) if $a,\tilde{a}\in\overline{H^{\infty}}+\mathcal{C}$ and $b,\tilde{b}\in H^{\infty}+\mathcal{C}$ (resp. $a,\tilde{a}\in\overline{H^{\infty}}+\mathcal{R}$ and $b,\tilde{b}\in H^{\infty}+\mathcal{R}$).
(iii) $\Sigma_{a,b}\eta I-\eta\Sigma_{a,b}$ is compact (resp. finite rank) if $\eta\in\mathcal{C}$ (resp. $\eta\in\mathcal{R}$).
In particular, it is clear from (i) and (ii) in Proposition 2.7 that $[S_{a,b},S_{\tilde{a},\tilde{b}}]=0$, that is, $S_{a,b}$ and $S_{\tilde{a},\tilde{b}}$ commute, if one of the following conditions hold:
$a=b\quad$ $\hbox{and}\quad\tilde{a}=\tilde{b};$ (2.5)
$a,\tilde{a}\in H^{\infty}\quad$ $\hbox{and}\quad b,\tilde{b}\in\overline{H^{\infty}};$ (2.6)
$a=\lambda\tilde{a}+\mu\quad$ $\hbox{and}\quad b=\lambda\tilde{b}+\mu,\quad\hbox{with}\quad\lambda,\mu\in\mathbb{C}.$ (2.7)
The converse was provided in [15, Thm. 4.1], and so we have:
Proposition 2.10.
$[S_{a,b},S_{\tilde{a},\tilde{b}}]=0$ if and only if one of the conditions (2.5)–(2.7) holds.
A similar property holds for $[\Sigma_{a,b},\Sigma_{\tilde{a},\tilde{b}}]$ by (2.2).
As an illustration of the results in Proposition 2.8 we have, for nondegenerate $(a,b)$, that $[S_{a,b},M_{z}]$ and $[S_{a,b},M_{\overline{z}}]$ are rank-one operators, such that
$\begin{array}[]{lcl}[S_{a,b},M_{z}](f)=&S_{a,b}zf-zS_{a,b}f=&(a-b)(\overline{z}\overline{f_{-}})(0),\\ {[S_{a,b},M_{\overline{z}}](f)}=&S_{a,b}\overline{z}f-\overline{z}S_{a,b}f=&(b-a)\overline{z}f_{+}(0),\end{array}$ (2.8)
where $f_{\pm}=P^{\pm}f$.
Analogously, $[\Sigma_{a,b},M_{z}]$ and $[\Sigma_{a,b},M_{\overline{z}}]$ are also rank-one operators, such that
$\begin{array}[]{lcl}[\Sigma_{a,b},M_{z}](f)=&\Sigma_{a,b}zf-z\Sigma_{a,b}f=&[zP^{-}(a-b)f](\infty),\\ {[\Sigma_{a,b},M_{\overline{z}}](f)}=&\Sigma_{a,b}\overline{z}f-\overline{z}\Sigma_{a,b}f=&-\overline{z}[P^{+}(a-b)f](0).\end{array}$ (2.9)
From Proposition 2.10 we also get the following.
Corollary 2.11.
If $(a,b)$ is nondegenerate and $\eta\in L^{\infty}$ then
$S_{a,b}\eta I=\eta S_{a,b}\iff\eta\in\mathbb{C}\iff\Sigma_{a,b}\eta I=\eta\Sigma_{a,b}.$
For $\eta\in L^{\infty}\setminus\mathbb{C}$ one may then ask which functions $f\in L^{2}$ satisfy the relation $S_{a,b}\eta f=\eta S_{a,b}f$. We have the following from [4, Prop. 3.4].
Proposition 2.12.
For $\eta\in L^{\infty}$, $(a,b)$ nondegenerate, and $f\in L^{2}$,
$\eta S_{a,b}f=S_{a,b}(\eta f)$ $\iff$ $f\in\mathop{\rm ker}\nolimits H_{\eta}\oplus\mathop{\rm ker}\nolimits\tilde{H}_{\eta}$
$\iff$ $\eta f_{+}\in H^{2}_{+},\eta f_{-}\in H^{2}_{-}$
$\iff$ $\eta f_{+}=P^{+}(\eta f)$
$\iff$ $\eta f_{-}=P^{-}(\eta f),$
where the Hankel operators $H_{\eta}$ and $\tilde{H}_{\eta}$ are defined by $H_{\eta}=P^{-}\eta P^{+}{}_{|H^{2}_{+}}$ and $\tilde{H}_{\eta}=P^{+}\eta P^{-}{}_{|H^{2}_{-}}$, and we write $f_{\pm}=P^{\pm}f$.
So
$\mathop{\rm ker}\nolimits[S_{a,b},\eta I]=\mathop{\rm ker}\nolimits H_{\eta}\oplus\mathop{\rm ker}\nolimits\tilde{H}_{\eta}.$
Thus, in particular, for nondegenerate $(a,b)$, the space of all functions $f\in L^{2}$ such that $S_{a,b}zf=zS_{a,b}f$ is given by
$\mathop{\rm ker}\nolimits[S_{a,b},M_{z}]=\overline{z}H^{2}_{-}\oplus H^{2}_{+}=\{f\in L^{2}:(\overline{z}\overline{f_{-}})(0)=0\}$ (2.10)
and, analogously,
$S_{a,b}\overline{z}f=\overline{z}S_{a,b}f\iff f\in H^{2}_{-}\oplus zH^{2}_{+}\iff f_{+}(0)=0.$ (2.11)
We may prove a similar result for transposed paired kernels.
Proposition 2.13.
For $\eta\in L^{\infty}$, $(a,b)$ nondegenerate, and $f\in L^{2}$, we have $\Sigma_{a,b}\eta f=\eta\Sigma_{a,b}f\iff(a-b)f\in\mathop{\rm ker}\nolimits[S_{a,b},\eta I]$, and hence $\mathop{\rm ker}\nolimits[\Sigma_{a,b},\eta I]=\{f\in L^{2}:(a-b)f\in\mathop{\rm ker}\nolimits H_{\eta}\oplus\mathop{\rm ker}\nolimits\tilde{H}_{\eta}\}$.
Proof.
$P^{+}a\eta f+P^{-}b\eta f=\eta P^{+}af+\eta P^{-}bf\iff P^{+}a\eta f-P^{+}b\eta f=\eta P^{+}af-\eta P^{+}bf$. This holds if and only if $P^{+}\eta(a-b)f=\eta P^{+}(a-b)f$, and so by Proposition 2.12 if and only if $(a-b)f\in\mathop{\rm ker}\nolimits[S_{a,b},\eta I]$. ∎
Finally, one may ask when a transposed paired operator $\Sigma_{a,b}=S_{\overline{a},\overline{b}}^{*}$ is also a paired operator. We have the following answer from [15, Prop. 2.2, Cor. 2.3], to which we provide an alternative proof.
Proposition 2.14.
$S^{*}_{a,b}=S_{c,d}$ for some $c,d\in L^{\infty}$ if and only if $a-b\in\mathbb{C}$, and, in that case, $S^{*}_{a,b}=S_{\overline{a},\overline{b}}$. Thus $S_{a,b}$ is self-adjoint if and only if $a$ and $b$ are real-valued functions with $a-b\in\mathbb{R}$.
Proof.
Suppose that
$P^{+}\overline{a}f+P^{-}\overline{b}f=cP^{+}f+dP^{-}f\qquad\hbox{for all}\quad f\in L^{2}.$ (2.12)
Then, for $f=1$, we get that $P^{+}\overline{a}+P^{-}\overline{b}=c$, so $P^{+}\overline{a}=P^{+}c$ and $P^{-}\overline{b}=P^{-}c$. On the other hand, taking $f=\overline{z}$, we get $P^{+}(\overline{a}\overline{z})+P^{-}(\overline{b}\overline{z})=d\overline{z}$, i.e.,
$\overline{z}(P^{+}\overline{a})-\overline{z}(P^{+}\overline{a})(0)+\overline{z}(P^{-}\overline{b})+\overline{z}(P^{+}\overline{b})(0)=d\overline{z},$
that is, $\overline{z}c-\overline{z}k=\overline{z}d$ with $k=(P^{+}(\overline{a}-\overline{b}))(0)\in\mathbb{C}$.
Therefore, $c=d+k$ and it follows from (2.12) that
$P^{+}(\overline{a}-k)I+P^{-}\overline{b}I=dI$
so, by Corollary 2.2, $\overline{a}-k=\overline{b}=d=c-k$; hence $a-b\in\mathbb{C}$ and $S^{*}_{a,b}=S_{c,d}$. The converse is straightforward. ∎
Turning our attention to kernels, and given the close connection between paired operators of the form $S_{a,b}$ and transposed paired operators of the form $\Sigma_{a,b}$, which was mentioned in the introduction and the beginning of this section, one might expect their kernels to have very similar properties. Moreover, we have the following simple relation, which generalizes [4, Prop. 6.4].
Proposition 2.15.
For nondegenerate $(a,b)$ there is a well-defined injective linear mapping $\mathcal{J}:\mathop{\rm ker}\nolimits\Sigma_{a,b}\to\mathop{\rm ker}\nolimits S_{a,b}$ given by
$\mathcal{J}\psi=(a-b)\psi.$ (2.13)
If there exist $a^{\prime},b^{\prime}\in L^{\infty}$ such that $aa^{\prime}+bb^{\prime}=1$, then $\mathcal{J}$ is an isomorphism with inverse
$\mathcal{J}^{-1}=a^{\prime}P^{-}-b^{\prime}P^{+}:\mathop{\rm ker}\nolimits S_{a,b}\to\mathop{\rm ker}\nolimits\Sigma_{a,b},$
and in that case
$(a-b)\mathop{\rm ker}\nolimits\Sigma_{a,b}=\mathop{\rm ker}\nolimits S_{a,b}.$
Proof.
For the last part, note that $a^{\prime}P^{-}(a-b)\psi-b^{\prime}P^{+}(a-b)\psi=a^{\prime}(a\psi)+b^{\prime}(b\psi)=\psi$ when $P^{+}(a\psi)=P^{-}(b\psi)=0$, i.e., when $\psi\in\mathop{\rm ker}\nolimits\Sigma_{a,b}$. ∎
Somewhat surprisingly, in spite of some similarities, the two types of kernels present significantly different and even contrasting properties, as shown in the next two sections, where we study each type of kernel separately.
3 Paired kernels
We start by noting that the kernels of paired operators, called paired kernels, are determined by the solutions of a Riemann–Hilbert problem for the unknowns $\phi_{\pm}=P^{\pm}\phi$:
$\phi\in\mathop{\rm ker}\nolimits S_{a,b}\iff a\phi_{+}+b\phi_{-}=0.$ (3.1)
It is clear from (3.1) that paired kernels are closely connected with Toeplitz kernels. Indeed, defining the Toeplitz operator
$T_{a/b}:\mathcal{D}_{a/b}:=\{f\in H^{2}_{+}:(a/b)f\in L^{2}\}\to H^{2}_{+},$ (3.2)
$f\mapsto P^{+}(a/b)f,$ (3.3)
which is bounded on $H^{2}_{+}$ if $a/b\in L^{\infty}$, we have from (3.3) that
$P^{+}\mathop{\rm ker}\nolimits S_{a,b}=\mathop{\rm ker}\nolimits T_{a/b}.$ (3.4)
Nevertheless, no nonzero paired kernel can be a Toeplitz kernel, or even be contained in $H^{2}_{+}$. This is a consequence of the following [8, Rem. 4.2].
Proposition 3.1.
If $\phi\in\mathop{\rm ker}\nolimits S_{a,b}$ then $\phi=0\iff\phi_{+}=0\iff\phi_{-}=0$. Therefore,
$\mathop{\rm ker}\nolimits S_{a,b}\cap H^{2}_{+}=\mathop{\rm ker}\nolimits S_{a,b}\cap H^{2}_{-}=\{0\}.$ (3.5)
Clearly, for any measurable complex-valued function $\eta$ defined a.e. on $\mathbb{T}$, such that $\eta\neq 0$ a.e., we have that
$a\phi_{+}+b\phi_{-}=0\iff a\eta\phi_{+}+b\eta\phi_{-}=0$
so, for any $\eta\in L^{\infty}$ with $\eta\neq 0$ a.e.,
$\mathop{\rm ker}\nolimits S_{a,b}=\mathop{\rm ker}\nolimits S_{a\eta,b\eta}.$ (3.6)
The next results show that this is the only case where two paired kernels are equal.
Proposition 3.2.
[4, Prop. 4.3] If $\mathop{\rm ker}\nolimits S_{a,b}\neq\{0\}$ then, for any nondegenerate $\tilde{a},\tilde{b}$ satisfying the assumption (1.12) we have
$\mathop{\rm ker}\nolimits S_{a,b}=\mathop{\rm ker}\nolimits S_{\tilde{a},\tilde{b}}\iff a\tilde{b}=\tilde{a}b.$ (3.7)
As a consequence, $\mathop{\rm ker}\nolimits S_{a,b}$ is uniquely determined by the quotient $a/b$, which is the same as the quotient $-\phi_{-}/\phi_{+}$ for any nonzero element of $\mathop{\rm ker}\nolimits S_{a,b}$, by (3.1). Thus we get the following striking property of paired kernels [4, Sec. 4].
Theorem 3.3.
For each $\phi\in L^{2}\setminus\{0\}$ there is one and only one paired kernel to which $\phi$ belongs. If $\phi_{+}=I_{+}O_{+}$ and $\phi_{-}=I_{-}O_{-}$, where $I_{+},\overline{I_{-}}$ are inner functions and $O_{\pm}$ outer in $H^{2}_{\pm}$, then
$\phi\in\mathop{\rm ker}\nolimits S_{a,b},\quad\hbox{with}\quad a=\overline{I_{+}}H_{1+}\overline{h_{2}+},\quad b=-z\overline{I_{-}}\,\overline{h_{1+}}H_{2+},$ (3.8)
where $H_{1+},H_{2+},h_{1+},h_{2+}\in H^{\infty}$ are such that
$O^{-1}_{+}=\frac{H_{1+}}{H_{2+}},\quad z\overline{O^{-1}_{-}}=\frac{h_{1+}}{h_{2+}}.$ (3.9)
Corollary 3.4.
If $\mathop{\rm ker}\nolimits S_{a.b}\cap\mathop{\rm ker}\nolimits S_{\tilde{a},\tilde{b}}\neq\{0\}$, then $\mathop{\rm ker}\nolimits S_{a,b}=\mathop{\rm ker}\nolimits S_{\tilde{a},\tilde{b}}$.
Corollary 3.5.
$\mathop{\rm ker}\nolimits S_{a,b}\subseteq\mathop{\rm ker}\nolimits S_{\tilde{a},\tilde{b}}\iff\mathop{\rm ker}\nolimits S_{a,b}=\{0\}$ or $\mathop{\rm ker}\nolimits S_{a,b}=\mathop{\rm ker}\nolimits S_{\tilde{a},\tilde{b}}$.
When do we have $\mathop{\rm ker}\nolimits S_{a,b}\neq\{0\}$? From the proof of [4, Thm. 4.6] we obtain a necessary and sufficient condition for a nonzero kernel.
Proposition 3.6.
$\mathop{\rm ker}\nolimits S_{a,b}\neq\{0\}$ if and only if
$\frac{a}{b}=I_{+}O_{-}O_{+}^{-1},$ (3.10)
where $I_{+}$ is inner and $O_{\pm}$ are outer in $H^{2}_{\pm}$.
Note that this coincides with the necessary and sufficient condition for $\mathop{\rm ker}\nolimits T_{a/b}$ to be nontrivial, when $a/b\in L^{\infty}$, obtained in [20, 7].
An analogue of Coburn’s lemma for Toeplitz kernels [10] also holds for paired kernels. To see this, however, it is useful to reformulate Coburn’s lemma, usually expressed as:
$\hbox{if}\quad g\neq 0\quad\hbox{then either}\quad\mathop{\rm ker}\nolimits T_{g}=\{0\}\quad\hbox{or}\quad\mathop{\rm ker}\nolimits T^{*}_{g}=\{0\}.$
Note that, while the adjoint of a Toeplitz operator is also a Toeplitz operator, the adjoint of a paired operator is not, in general, a paired operator of the same type. Noting, however, that we can always assume that the symbol in $\mathop{\rm ker}\nolimits T_{g}$ satisfies $|g|=1$ [16] and, on the other hand,
$\mathop{\rm ker}\nolimits T_{g}\cong\mathop{\rm ker}\nolimits S_{g,1},\quad\mathop{\rm ker}\nolimits T^{*}_{g}=\mathop{\rm ker}\nolimits T_{\overline{g}}\cong\mathop{\rm ker}\nolimits S_{\overline{g},1}=\mathop{\rm ker}\nolimits S_{1,g}$
by (1.11), we can also formulate Coburn’s lemma, equivalently, as follows, in terms of paired operators.
Theorem 3.7.
(Coburn’s lemma) If $|g|=1$ then either $\mathop{\rm ker}\nolimits S_{g,1}=\{0\}$ or $\mathop{\rm ker}\nolimits S_{1,g}=\{0\}$.
We can formulate the following analogues for paired kernels, which slightly generalize Theorem 6.8 in [4].
Theorem 3.8.
(i) Either $\mathop{\rm ker}\nolimits S_{a,b}=\{0\}$ or $\mathop{\rm ker}\nolimits S_{b,a}=\{0\}$.
(ii) If there exist $a^{\prime},b^{\prime}\in L^{\infty}$ such that $aa^{\prime}+bb^{\prime}=1$, then either $\mathop{\rm ker}\nolimits S_{a,b}=\{0\}$ or $\mathop{\rm ker}\nolimits S^{*}_{a,b}=\{0\}$.
4 Transposed paired kernels
It is not difficult to see that
$\mathop{\rm ker}\nolimits\Sigma_{a,b}$ $=$ $\{\phi\in L^{2}:P^{+}a\phi=P^{-}b\phi=0\}$ (4.1)
$=$ $\{\phi\in L^{2}:a\phi=\psi_{-},\quad b\phi=\psi_{+}\quad\hbox{with}\quad\psi_{\pm}\in H^{2}_{\pm}\}$ (4.2)
$=$ $L^{2}\cap a^{-1}H^{2}_{-}\cap b^{-1}H^{2}_{+},\quad\hbox{for non-degenerate }(a,b).$ (4.3)
Thus, in contrast with paired kernels, the elements of a transposed paired kernel are not, in general, determined by the solution of a scalar Riemann–Hilbert problem as in (3.1). As a consequence, transposed paired kernels possess properties that are different, and even in contrast, with those of paired kernels and they have different relations with Toeplitz kernels.
Defining $T_{a/b}$ as in (3.2) and (3.3), and taking into account the fact that
$a^{-1}H^{2}_{-}\cap b^{-1}H^{2}_{+}=b^{-1}\left(\frac{b}{a}H^{2}_{-}\cap H^{2}_{+}\right),$
we have, from (4.3):
Proposition 4.1.
(i) $\mathop{\rm ker}\nolimits\Sigma_{a,b}=\left(\dfrac{1}{b}\mathop{\rm ker}\nolimits T_{a/b}\right)\cap L^{2}$;
$b\mathop{\rm ker}\nolimits\Sigma_{a,b}\subseteq\mathop{\rm ker}\nolimits T_{a/b}$ (4.4)
and, if $b\in\mathcal{G}L^{\infty}$, then
$b\mathop{\rm ker}\nolimits\Sigma_{a,b}=\mathop{\rm ker}\nolimits T_{a/b}.$ (4.5)
In particular, we see from (4.5) that
$b\in\mathcal{G}H^{\infty}\implies\mathop{\rm ker}\nolimits\Sigma_{a,b}=\frac{1}{b}\mathop{\rm ker}\nolimits T_{a/b}=\mathop{\rm ker}\nolimits T_{a},$ (4.6)
so, in contrast with the paired kernels studied in Section 3, nontrivial transposed paired kernels can be Toeplitz kernels. Indeed, we have the following.
Proposition 4.2.
Every Toeplitz kernel can be written as
$\mathop{\rm ker}\nolimits T_{a}=\mathop{\rm ker}\nolimits\Sigma_{a,b}\qquad\hbox{for any}\quad b\in H^{\infty},\quad\hbox{outer}.$ (4.7)
Proof.
If $b\in H^{\infty}$ is outer, then $b^{-1}H^{2}_{+}\cap L^{2}=H^{2}_{+}$ and it follows from (4.3) that $\mathop{\rm ker}\nolimits\Sigma_{a,b}=a^{-1}H^{2}_{-}\cap H^{2}_{+}=\mathop{\rm ker}\nolimits T_{a}$. ∎
Corollary 4.3.
If $a\in\overline{H^{\infty}}$ and $b\in H^{\infty}$, with $b$ outer, then $\mathop{\rm ker}\nolimits S_{a,b}$ is a model space $K_{\theta}$, where $\theta$ is the inner factor of $\overline{a}$. In particular,
$K_{\theta}=\mathop{\rm ker}\nolimits\Sigma_{\overline{\theta},1}.$ (4.8)
These results naturally raise the question of how to characterise $\mathop{\rm ker}\nolimits\Sigma_{a,b}\cap H^{2}_{\pm}$, and, in particular, when $\mathop{\rm ker}\nolimits\Sigma_{a,b}$ is contained in $H^{2}_{\pm}$ or vice-versa. The answer reveals some interesting connections of transposed paired kernels with Toeplitz and Hankel operators.
Proposition 4.4.
$\mathop{\rm ker}\nolimits\Sigma_{a,b}\cap H^{2}_{+}=\mathop{\rm ker}\nolimits T_{a}\cap\mathop{\rm ker}\nolimits H_{b},$ (4.9)
$\mathop{\rm ker}\nolimits\Sigma_{a,b}\cap H^{2}_{-}=\mathop{\rm ker}\nolimits\widetilde{H}_{a}\cap\mathop{\rm ker}\nolimits\widetilde{T}_{b}.$
Proof.
(i) and (ii) follow straightforwardly from (4.1). ∎
The following lemma will be used several times in what follows.
Lemma 4.5.
If $\mathop{\rm ker}\nolimits\Sigma_{a,b}\neq\{0\}$ then there exist $\phi\in\mathop{\rm ker}\nolimits\Sigma_{a,b}$ such that $a\phi=\psi_{-}$ with $\psi_{-}\in H^{2}_{-}$, outer, and $\check{\phi}\in\mathop{\rm ker}\nolimits\Sigma_{a,b}$ such that $b\check{\phi}=\psi_{+}$ with $\psi_{+}\in H^{2}_{+}$, outer.
Proof.
Let $0\neq\tilde{\phi}\in\mathop{\rm ker}\nolimits\Sigma_{a,b}$; then $a\tilde{\phi}=\psi_{-}$ and $b\tilde{\phi}=\tilde{\psi}_{+}$ where $\tilde{\psi}_{\pm}\in H^{2}_{\pm}$. Let $\tilde{\psi}_{-}=I_{-}\psi_{-}$, where $\overline{I_{-}}$ is inner and $\psi_{-}$ is outer in $H^{2}_{-}$. Then we have $a(\overline{I_{-}}\tilde{\phi})=\psi_{-}$, $b(\overline{I_{-}}\tilde{\phi})=\overline{I_{-}}\tilde{\psi}_{+}\in H^{2}_{+}$, so $\phi:=\overline{I_{-}}\tilde{\phi}$ satisfies $a\phi=\psi_{-}$ with $\psi_{-}\in H^{2}_{-}$, outer. We can prove similarly the second part of the proposition. ∎
Proposition 4.6.
$\mathop{\rm ker}\nolimits\Sigma_{a,b}\subseteq H^{2}_{+}\iff\mathop{\rm ker}\nolimits\Sigma_{a,b}=\mathop{\rm ker}\nolimits T_{a}\cap\mathop{\rm ker}\nolimits H_{b}$, and this occurs if and only if there exists $\psi^{o}_{+}\in H^{2}_{+}$, outer, such that $\psi^{o}_{+}/b\in H^{2}_{+}$.
Proof.
The first equivalence is a consequence of Proposition 4.4 (i). For the second equivalence, suppose that $\psi_{+}^{o}/b\in H^{2}_{+}$ with $\psi^{o}_{+}\in H^{2}_{+}$, outer. Then for any $\phi\in\mathop{\rm ker}\nolimits\Sigma_{a,b}$ we have that $b\phi=\psi_{+}\in H^{2}_{+}$, so
$\phi=\frac{\psi_{+}}{b}=\frac{\psi_{+}(\psi_{+}^{o}/b)}{\psi_{+}^{o}}\in\mathcal{N}^{+}\cap L^{2}=H^{2}_{+}.$
(Here $\mathcal{N}_{+}$ is the Smirnoff class [13, Thm. 2.11].)
Conversely, if $\mathop{\rm ker}\nolimits\Sigma_{a,b}\subseteq H^{2}_{+}$, then by Lemma 4.5 there exists $\psi^{o}_{+}\in H^{2}_{+}$, outer, such that $b\check{\phi}=\psi^{o}_{+}$ and $\check{\phi}\in\mathop{\rm ker}\nolimits\Sigma_{a,b}\subseteq H^{2}_{+}$, so $\psi^{o}_{+}/b\in H^{2}_{+}$. ∎
Proposition 4.7.
$H^{2}_{+}\subseteq\mathop{\rm ker}\nolimits\Sigma_{a,b}\iff a=0\quad\hbox{and}\quad b\in H^{\infty}$ (4.10)
and
$H^{2}_{+}=\mathop{\rm ker}\nolimits\Sigma_{a,b}\iff a=0,\ b\in H^{\infty}\quad\hbox{with}\quad b\hbox{ outer}.$
Proof.
If $H^{2}_{+}\subseteq\mathop{\rm ker}\nolimits\Sigma_{a,b}$ then $H^{2}_{+}=\mathop{\rm ker}\nolimits\Sigma_{a,b}\cap H^{2}_{+}=\mathop{\rm ker}\nolimits T_{a}\cap\mathop{\rm ker}\nolimits H_{b}$. It follows that $\mathop{\rm ker}\nolimits T_{a}=H^{2}_{+}$, so $a=0$ and $\mathop{\rm ker}\nolimits H_{b}=H^{2}_{+}$, so $b\in H^{\infty}$. The converse is obvious. The conditions for equality follow from (4.10) and Proposition 4.6. ∎
Analogous results hold if we replace $H^{2}_{+}$ with $H^{2}_{-}$, showing that we can also have $\mathop{\rm ker}\nolimits\Sigma_{a,b}\subseteq H^{2}_{-}$, $H^{2}_{-}\subseteq\mathop{\rm ker}\nolimits\Sigma_{a,b}$ and $\mathop{\rm ker}\nolimits\Sigma_{a,b}=H^{2}_{-}$. In particular, if $h_{+}\in H^{\infty}$ is outer, then
$\mathop{\rm ker}\nolimits\Sigma_{\overline{h_{+}},\theta}=\overline{z}\overline{K_{\theta}}=\overline{\theta}K_{\theta}\subseteq H^{2}_{-}.$ (4.11)
Now, it is clear that, in general, for $\eta\in L^{\infty}$,
$\mathop{\rm ker}\nolimits\Sigma_{a,b}\neq\mathop{\rm ker}\nolimits\Sigma_{a\eta,b\eta},$ (4.12)
in contrast with (3.6). The question then is when two transposed kernels are equal, or related by inclusion. Here, again, the results are in contrast with those previously obtained for paired kernels. First, however, we prove the following.
Proposition 4.8.
If $\phi\in\mathop{\rm ker}\nolimits\Sigma_{a,b}$ and $\phi\neq 0$, then $\phi$ cannot vanish on a subset of $\mathbb{T}$ with positive measure.
Proof.
If $\phi$ vanishes on a set $E\subset\mathbb{T}$ with poisitve measure, then $a\phi=\psi_{-}=0$ (since it is in $H^{2}_{-}$ and vanishes on $E$), and $b\phi=\psi_{+}=0$ (similarly), so $\phi=0$. ∎
Proposition 4.9.
Let $\mathop{\rm ker}\nolimits\Sigma_{a,b}\neq\{0\}$. Then:
(i) $\mathop{\rm ker}\nolimits\Sigma_{a,b}\subseteq\mathop{\rm ker}\nolimits\Sigma_{\tilde{a},\tilde{b}}$ if and only if $\tilde{a}/a=\tilde{\psi}_{-}/\psi_{-}$ with $\psi_{-},\tilde{\psi}_{-}\in H^{2}_{-}$ and $\psi_{-}$ outer, and $\tilde{b}/b=\tilde{\psi}_{+}/\psi_{+}$ with $\psi_{+},\tilde{\psi}_{+}\in H^{2}_{+}$ and $\psi_{+}$ outer;
(ii) $\mathop{\rm ker}\nolimits\Sigma_{a,b}=\mathop{\rm ker}\nolimits\Sigma_{\tilde{a},\tilde{b}}$ if and only if $\tilde{a}/a$ and $\tilde{b}/b$ can be written as in (i) with $\tilde{\psi}_{-}$ and $\tilde{\psi}_{+}$ outer in $H^{2}_{-}$ and $H^{2}_{+}$, respectively;
(iii) if $h_{-}\in\overline{H^{\infty}}$ and $h_{+}\in H^{\infty}$, then
$\mathop{\rm ker}\nolimits\Sigma_{a,b}\subseteq\mathop{\rm ker}\nolimits\Sigma_{ah_{-},bh_{+}},$
where the inclusion is strict if $\overline{h_{-}}$ or $h_{+}$ has a nonconstant inner factor, and equality if $h_{-}$ and $h_{+}$ are outer functions.
Proof.
(i) Let $\{0\}\neq\mathop{\rm ker}\nolimits\Sigma_{a,b}\subseteq\mathop{\rm ker}\nolimits\Sigma_{\tilde{a},\tilde{b}}$ and let $\phi\in\mathop{\rm ker}\nolimits\Sigma_{a,b}$ with $a\phi=\psi_{-}\in H^{2}_{-}$, $\psi_{-}$ outer (see Lemma 4.5). Then we have
$a\phi=\psi_{-},\quad\tilde{a}\phi=\tilde{\psi}_{-}\in H^{2}_{-},$
so $\tilde{a}/a=\tilde{\psi}_{-}/\psi_{-}$ with $\psi_{-},\tilde{\psi}_{-}\in H^{2}_{-}$ and $\psi_{-}$ outer. Analogously, using Lemma 4.5, we can write $\tilde{b}/b=\tilde{\psi}_{+}/\psi_{+}$ with $\psi_{+},\tilde{\psi}_{+}\in H^{2}_{+}$ and $\psi_{+}$ outer.
Conversely, if $\tilde{a}/a=\tilde{\psi}_{-}/\psi_{-}$ and $\tilde{b}/b=\tilde{\psi}_{+}/\psi_{+}$, where $\tilde{\psi}_{\pm},\psi_{\pm}\in H^{2}_{\pm}$, with $\psi_{\pm}$, and $\phi\in\mathop{\rm ker}\nolimits\Sigma_{a,b}$, then $a\phi=\phi_{-}\in H^{2}_{-}$, $b\phi=\phi_{+}\in H^{2}_{+}$, and
$\tilde{a}\phi$ $=$ $\frac{\tilde{a}}{a}(a\phi)=\frac{\tilde{\psi}_{-}}{\psi_{i}}\in\overline{z}\overline{\mathcal{N}_{+}}\cap L^{2}=H^{2}_{-},$
$\tilde{b}\phi$ $=$ $\frac{\tilde{b}}{b}(b\phi)=\frac{\tilde{\psi}_{+}}{\psi_{+}}\phi_{+}\in\mathcal{N}_{+}\cap L^{2}=H^{2}_{+}$
where $\mathcal{N}_{+}$ is the Smirnoff class [13, Thm. 2.11]. So $\phi\in\mathop{\rm ker}\nolimits\Sigma_{\tilde{a},\tilde{b}}$.
(ii) If $\psi_{-},\tilde{\psi}_{-}\in H^{2}_{-}$ are outer and $\psi_{+},\tilde{\psi}_{+}\in H^{2}_{+}$ are outer, then by (i) we have $\mathop{\rm ker}\nolimits\Sigma_{a,b}\subseteq\mathop{\rm ker}\nolimits\Sigma_{\tilde{a},\tilde{b}}$ and $\mathop{\rm ker}\nolimits\Sigma_{\tilde{a},\tilde{b}}\subseteq\mathop{\rm ker}\nolimits\Sigma_{a,b}$.
Conversely, assume that $\tilde{a}/a=\tilde{\psi}_{-}/\psi_{-}$ and $\tilde{b}/b=\tilde{\psi}_{+}/\psi_{+}$ as in (i), where, for instance $\tilde{\psi}_{-}\in H^{2}_{-}$ is not outer, so $\tilde{a}/a=\overline{\theta}\eta_{-}/\psi_{-}$ with $\theta$ inner and $\eta_{-},\psi_{-}\in H^{2}_{-}$ outer. Then by (i) we have that $\mathop{\rm ker}\nolimits\Sigma_{a,b}\subseteq\mathop{\rm ker}\nolimits\Sigma_{\tilde{a},\tilde{b}}$, but we cannot have $\mathop{\rm ker}\nolimits\Sigma_{\tilde{a},\tilde{b}}\subseteq\mathop{\rm ker}\nolimits\Sigma_{a,b}$. Indeed, let $\psi\in\mathop{\rm ker}\nolimits\Sigma_{\tilde{a},\tilde{b}}$ be such that (see Lemma 4.5) $\tilde{a}\phi=\phi_{-}$ with $\phi_{-}$ outer in $H^{2}_{-}$. If $\phi\in\mathop{\rm ker}\nolimits\Sigma_{a,b}$ then we also have $a\phi=\chi_{-}\in H^{2}_{-}$; from this it follows that
$\tilde{a}\phi=a\overline{\theta}\frac{\eta_{-}}{\psi_{-}}\phi=\phi_{-}\quad\hbox{and}\quad a\phi=\chi_{-},$
so that $\phi_{-}=\overline{\theta}\dfrac{\eta_{-}}{\psi_{-}}\chi_{-}$, which is impossible because $\phi_{-}$ is outer in $H^{2}_{-}$.
(iii) This is a consequence of (i) and (ii). ∎
When is $\mathop{\rm ker}\nolimits\Sigma_{a,b}\neq\{0\}$? We have the following necessary and sufficient condition.
Proposition 4.10.
$\mathop{\rm ker}\nolimits\Sigma_{a,b}\neq\{0\}$ if and only if there exist $I_{-},O_{-},O_{+}$ such that $\overline{I_{-}}$ is inner, $O_{\pm}\in H^{2}_{\pm}$ are outer and
$\frac{a}{b}=I_{-}O_{-}O^{-1}_{+},\quad\frac{O_{-}}{a},\frac{O_{+}}{b}\in L^{2}.$ (4.13)
Proof.
Let $\phi\in\mathop{\rm ker}\nolimits\Sigma_{a,b}$, $\phi\neq 0$. Since $a\phi=\psi_{-}\in H^{2}_{-}$ and $b\phi=\psi_{+}\in H^{2}_{+}$, we have that $a/b=\psi_{-}/\psi_{+}$ can be written as in (4.13), where $\psi_{-}=J_{-}O_{-}$ and $\psi_{+}=J_{+}O_{+}$ with $\overline{J_{-}},J_{+}$ inner and $O_{\pm}\in H^{2}_{\pm}$ outer, and then $I_{-}=J_{-}/J_{+}$.
We then have $O_{-}/a=\overline{J_{-}}\phi\in L^{2}$ and $O_{+}/b=\overline{J_{+}}\phi\in L^{2}$.
Conversely, if (4.13) holds, then
$a\frac{O_{+}}{b}$ $=$ $(I_{-}O_{-}O^{-1}_{+}b)\frac{O_{+}}{b}=I_{-}O_{-}\in H^{2}_{-},$
$b\frac{O_{+}}{b}$ $=$ $O_{+}\in H^{2}_{+},$
so $0\neq O_{+}/b\in\mathop{\rm ker}\nolimits\Sigma_{a,b}$. ∎
Remark 4.11.
Comparing (4.13) with (3.10) in Proposition 3.6, and taking Proposition 2.15 into account, we see that if $\mathop{\rm ker}\nolimits\Sigma_{a,b}\neq\{0\}$ then $\mathop{\rm ker}\nolimits S_{a,b}\neq\{0\}$. However, we may have $\mathop{\rm ker}\nolimits\Sigma_{a,b}=\{0\}$ and $\mathop{\rm ker}\nolimits S_{a,b}\neq\{0\}$. For example, let $a(z)=1+1/z$ and $b(z)=z+1$. Then with $f(z)=1-1/z$ we have $f\in\mathop{\rm ker}\nolimits S_{a,b}$ since
$(1+1/z)P^{+}f+(z+1)P^{-}f=1+1/z-(z+1)/z=0.$
But $\mathop{\rm ker}\nolimits\Sigma_{a,b}=\{0\}$, since if $P_{+}(af)=P_{-}(bf)=0$ then $(1+1/z)f\in H^{2}_{-}$ and $(z+1)f\in H^{2}_{+}$, which implies that $(z+1)f\in zH^{2}_{-}\cap H^{2}_{+}=\mathbb{C}$. But this is impossible unless $f=0$, since $f\in L^{2}(\mathbb{T})$.
Indeed, in this case $a/b=\overline{z}$ (i.e. $I_{-}=\overline{z}$, $O_{-}=O_{+}=1$), but $1/a,1/b\not\in L^{2}$.
Another striking difference between paired kernels and transposed paired kernels is expressed by the following property.
Proposition 4.12.
Not every $\phi\in L^{2}$ belongs to a transposed paired kernel. In particular, if $\phi$ vanishes on a set of positive measure in $\mathbb{T}$, then $\phi$ does not belong to $\mathop{\rm ker}\nolimits\Sigma_{a,b}$ for any $a,b$, unless $\phi=0$.
Proof.
If $\phi\in\mathop{\rm ker}\nolimits\Sigma_{a,b}$ then $P^{+}a\phi=P^{-}b\phi=0$. Thus $\phi=h_{+}/b$ for some $h_{+}\in H^{2}_{+}$ and, if $\phi$ vanishes on a set of positive measure, then $h_{+}=0$. ∎
However we do get an analogue of Coburn’s lemma in the case of transposed paired kernels, as in the case of paired kernels.
Proposition 4.13.
If $a,b\neq 0$ a.e., then either $\mathop{\rm ker}\nolimits\Sigma_{a,b}$ or $\mathop{\rm ker}\nolimits\Sigma_{b,a}$ is 00.
Proof.
This follows from Proposition 2.15 and Theorem 3.8. ∎
5 Invariance properties of paired kernels
It is easy to see that if $X$ and $T$ are commuting operators on a Banach space $\mathcal{X}$ then $\mathop{\rm ker}\nolimits T$ is invariant for $X$. Thus from the commutation relations of Section 2 (namely Proposition 2.10) we have the following.
Proposition 5.1.
[4, Prop. 5.1] If $a,\tilde{a}\in H^{\infty}$ and $b,\tilde{b}\in\overline{H^{\infty}}$, then $\mathop{\rm ker}\nolimits S_{a,b}$ is invariant for $S_{\tilde{a},\tilde{b}}$. If $a,\tilde{a}\in\overline{H^{\infty}}$ and $b,\tilde{b}\in H^{\infty}$, then $\mathop{\rm ker}\nolimits\Sigma_{a,b}$ is invariant for $\Sigma_{\tilde{a},\tilde{b}}$.
In particular, since for any inner function $\theta$, we have $\mathop{\rm ker}\nolimits\Sigma_{\overline{\theta},1}=K_{\theta}$ and, on the other hand, $\Sigma_{\overline{\theta},1}\Sigma_{\tilde{a},\tilde{b}}=\Sigma_{\tilde{a},\tilde{b}}\Sigma_{\overline{\theta},1}$ if $\tilde{a}\in\overline{H^{\infty}}$ and $\tilde{b}\in H^{\infty}$, we have the following generalization of the well-known $S^{*}$-invariance property of model spaces.
Corollary 5.2.
Model spaces are invariant subspaces for the transposed paired operators $\Sigma_{\tilde{a},\tilde{b}}$ with $\tilde{a}\in\overline{H^{\infty}}$ and $\tilde{b}\in H^{\infty}$.
Together with the notion of invariant subspaces for an operator $T$, one can define also the concept of almost-invariant subspaces for $T$.
Definition 5.3.
Let $\mathcal{H}$ be a Hilbert space, $X\in\mathcal{L}(\mathcal{H})$, and let $\mathcal{M}$ be a subspace of $\mathcal{H}$. We say that $\mathcal{M}$ is almost invariant with defect $m$ for $X$ if and only if
$X\mathcal{M}\subset\mathcal{M}\oplus\mathcal{F},$ (5.1)
where $\mathcal{F}$ is a finite-dimensional subspace of minimal dimension $m$.
It was shown by Popov and Tcaciuc [26] that every Hilbert space operator admits an almost-invariant half-space (that is, a subspace of infinite dimension and infinite codimension) of defect 1. The following result from [5] also holds.
Proposition 5.4.
Let $X,T\in\mathcal{L}(\mathcal{H})$, where $\mathcal{H}$ is a Hilbert space, such that $XT-TX$ is a finite-rank operator with rank $r$. Then $\mathop{\rm ker}\nolimits T$ is almost invariant with defect $m\leq r$ for $X$, i.e.,
$f\in\mathop{\rm ker}\nolimits T\implies Xf\in\mathop{\rm ker}\nolimits T\oplus\mathcal{F},$ (5.2)
where $\mathcal{F}$ is a subspace of dimension at most $r$.
From this proposition and Propositions 2.72.9 we then have the following.
Proposition 5.5.
If $a,\tilde{a}\in H^{\infty}+\mathcal{R}$ and $b,\tilde{b}\in\overline{H^{\infty}}+\mathcal{R}$, then $\mathop{\rm ker}\nolimits S_{a,b}$ is almost invariant for $S_{\tilde{a},\tilde{b}}$ with defect less than or equal to $\mathop{\rm rank}\nolimits[S_{a,b},S_{\tilde{a},\tilde{b}}]$.
A similar property holds for transposed paired operators. From (2.8) and (2.9) we have then:
Corollary 5.6.
$\mathop{\rm ker}\nolimits S_{a,b}$ and $\Sigma_{a,b}$ are almost invariant for multiplication by $z$ and $\overline{z}$ with defect 1.
Besides invariant and almost-invariant subspaces for an operator $T$, one can also define nearly-invariant subspaces. However, while the former are defined in a general Banach space setting, the notion of nearly $S^{*}$-invariant subspaces appeared in the particular context of $H^{2}_{+}$ in the work of Hitt, Hayashi and Sarason [17, 18, 27], motivated by questions of invariant subspaces and kernels of Toeplitz operators.
Definition 5.7.
Let $\mathcal{M}$ be a proper subspace of $H^{2}_{+}$. Then we say that $\mathcal{M}$ is nearly $S^{*}$-invariant if and only if
$f\in\mathcal{M},f(0)=0\implies S^{*}f\in\mathcal{M}.$ (5.3)
It is well known that kernels of Toeplitz operators are nearly $S^{*}$- invariant. However, kernels of Hankel operators $H_{g}:H^{2}_{+}\to H^{2}_{-}$ with $H_{g}\phi=P^{-}g\phi$ and truncated Toeplitz operators $A_{g}^{\theta}:K_{\theta}\to K_{\theta}$ with $A_{g}^{\theta}\phi=P_{\theta}g\phi$, where $P_{\theta}$ is the orthogonal projection onto $K_{\theta}$, are not in general nearly $S^{*}$-invariant, and the following weaker notion is of interest in those cases.
Definition 5.8.
[9] Let $\mathcal{M}$ be a proper subspace of $H^{2}_{+}$. Then we say that $\mathcal{M}$ is nearly $S^{*}$-invariant with defect $m$ if and only if
$f\in\mathcal{M},f(0)=0\implies S^{*}f\in\mathcal{M}\oplus\mathcal{F},$ (5.4)
where $\mathcal{F}$ is a finite-dimensional subspace of minimal dimension $m$.
A description of nearly $S^{*}$-invariant subspaces of defect $m$ was given in [9]. A nontrivial kernel of a Hankel operator, being of the form $\theta H^{2}_{+}$ for $\theta$ inner, is nearly $S^{*}$-invariant with defect 1 if $\theta(0)=0$, since if $f\in\theta H^{2}_{+}$ with $f(0)=0$ then $\bar{z}f\in\theta H^{2}+\mathbb{C}\bar{z}\theta$; while if $\theta(0)\neq 0$ it is already nearly $S^{*}$-invariant. Likewise the kernel of an asymmetric truncated Toeplitz operator is nearly $S^{*}$-invariant with defect 1, as shown in [25]; see also [5]. Similar observations apply to kernels of certain finite-rank perturbations of Toeplitz operators [19, 12].
Near $S^{*}$-invariance, with or without positive defect, can be seen and interpreted from various perspectives, leading to different generalizations of the notion of near invariance, allowing it to apply beyond the context of $H^{2}_{+}$. First, we see that (5.3) is equivalent to the condition
$f\in\mathcal{M},\overline{z}f\in H^{2}_{+}\implies\overline{z}f\in\mathcal{M}.$ (5.5)
So a natural generalization of Definition 5.8 is the following.
Definition 5.9.
[6] Let $\mathcal{M}$ be a proper subspace of $H^{2}_{+}$ and let $\eta$ be a complex-valued function defined a.e. on $\mathbb{T}$. We say that $\mathcal{M}$ is nearly $\eta$-invariant (or nearly $M_{\eta}$-invariant) if and only if
$f\in\mathcal{M},\ \eta f\in H^{2}_{+}\implies\eta f\in\mathcal{M},$ (5.6)
and, if $\eta\in L^{\infty}$, we also say that $\mathcal{M}$ is nearly $T_{\eta}$-invariant, since (5.6) is equivalent to
$f\in\mathcal{M},\ \eta f\in H^{2}_{+}\implies T_{\eta}f\in\mathcal{M}.$
Further, if
$f\in\mathcal{M},\ \eta f\in H^{2}_{+}\implies\eta f\in\mathcal{M}\oplus\mathcal{F},$ (5.7)
where $\mathcal{F}$ is a finite-dimensional subspace of minimal dimension $m$, then we say that $\mathcal{M}$ is nearly $\eta$-invariant with defect $m$.
It was shown in [6] that Toeplitz kernels are not only nearly $\overline{z}$-invariant (that is, nearly $S^{*}$-invariant) subspaces of $H^{2}_{+}$, but also nearly $\eta$-invariant for a wide class of functions $\eta$, including all functions in $\overline{H^{\infty}}$ and all rational functions without poles in the exterior of the unit disc. This property of near $\eta$-invariance provides important information characterizing Toeplitz kernels. For instance, it implies that a nontrivial Toeplitz kernel cannot be of the form $\alpha K$ with $\alpha$ inner and nonconstant and $K\subset H^{2}_{+}$, since Toeplitz kernels are nearly $\overline{\alpha}$-invariant. Nor can it be of the form $(z+a)K$ with $a\in\mathbb{D}\cup\mathbb{T}$ and $K\subset H^{2}_{+}$, since Toeplitz kernels are also nearly $1/(z+a)$-invariant.
To extend this notion beyond the particular setting of $H^{2}_{+}$, the following was introduced recently in [5].
Definition 5.10.
Let $\mathcal{H}$ be a Hilbert space and let $\mathcal{M}$, $\mathcal{N}$ be closed subspaces of $\mathcal{H}$, with $\mathcal{M}\subset\mathcal{N}$, such that for $X\in\mathcal{L}(\mathcal{H})$ we have
$f\in\mathcal{M},\ Xf\in\mathcal{N}\implies Xf\in\mathcal{M}.$ (5.8)
Then we say that $\mathcal{M}$ is nearly $X$-invariant with respect to $\mathcal{N}$ (for short, w.r.t. $\mathcal{N}$). Also, $\mathcal{M}$ is said to be nearly $X$-invariant w.r.t. $\mathcal{N}$ with defect $m$ if
$f\in\mathcal{M},\ Xf\in\mathcal{N}\implies Xf\in\mathcal{M}\oplus\mathcal{F},$ (5.9)
with $\mathcal{F}$ as above.
Note that, if $\mathcal{N}$ is a closed subspace and $P_{\mathcal{N}}$ denotes the orthogonal projection onto $\mathcal{N}$, then (5.8) is equivalent to
$f\in\mathcal{M},\ Xf\in\mathcal{N}\implies P_{\mathcal{N}}Xf\in\mathcal{M},$
so we also say that $\mathcal{M}$ is nearly $P_{\mathcal{N}}X_{|\mathcal{N}}$-invariant.
Taking, for instance $X=M_{\overline{z}}$, $\mathcal{N}=H^{2}_{+}$, $\mathcal{H}=L^{2}(\mathbb{T})$ and $P_{\mathcal{N}}X=S^{*}$, we recover (5.5) from (5.8).
There are obvious relations between the notions of almost-invariant and nearly-invariant subspaces for an operator, especially in the context of $H^{2}_{+}$. If $Xf\in\mathcal{N}$ for all $f\in\mathcal{M}$, then the space $\mathcal{M}$ is nearly $X$-invariant with defect $m$ w.r.t. $\mathcal{N}$ if and only if it is almost invariant with defect $m$ for $X$. This happens in particular if $\mathcal{N}=\mathcal{H}$. Moreover, as pointed out in [9], we also have that if $\mathcal{M}\subset H^{2}_{+}$ is almost invariant for $S^{*}$ with defect at most $m$, then it is nearly $S^{*}$-invariant with defect at most $m$. The converse is true if (i) all elements of $\mathcal{M}$ vanish at 00, or (ii) there exists $f_{0}\in\mathcal{M}$ with $f_{0}(0)\neq 0$ such that $S^{*}f_{0}\in\mathcal{M}+\mathcal{F}$.
Nevertheless, even when the notion of nearly $X$-invariant subspace of $\mathcal{H}$, with or without defect, coincides with that of almost-invariant subspace with a certain defect for the operator $X$, the two concepts express different points of view and address different questions on how the operator acts on a certain space.
For instance, in the context of $H^{2}_{+}$, saying that a subspace $\mathcal{M}$ is almost invariant with defect $m$ for an operator $X$ can be seen as answering the question of where $X(\mathcal{M})$ lies, if it is not contained in $\mathcal{M}$. Saying that $\mathcal{M}$ is nearly $X$-invariant, with or without defect, can be seen as answering the question of which elements of $\mathcal{M}$ are mapped into $\mathcal{M}$ (or $\mathcal{M}\oplus\mathcal{F}$) by $X$. Thus, considering the case of Toeplitz kernels, which are $S^{*}$-invariant if and only if the symbol is in $\overline{H^{\infty}}$, we may ask which elements $f\in\mathop{\rm ker}\nolimits T_{g}$ are mapped into $\mathop{\rm ker}\nolimits T_{g}$ by multiplication by $\overline{z}$, or by $S^{*}$, when $g\not\in\overline{H^{\infty}}$. The answer is given by $\overline{z}f\in H^{2}_{+}$, i.e.,
$f\in\mathop{\rm ker}\nolimits T_{g},\ \overline{z}f\in H^{2}_{+}\implies\overline{z}f\in\mathop{\rm ker}\nolimits T_{g},$ (5.10)
which is precisely the same as saying that $\mathop{\rm ker}\nolimits T_{g}$ is nearly $S^{*}$-invariant.
If we now consider instead the problem of how “far” $S^{*}(\mathop{\rm ker}\nolimits T_{g})$ is from $\mathop{\rm ker}\nolimits T_{g}$ when $g\not\in\overline{H^{\infty}}$, in the context of $H^{2}_{+}$, or the same question for $\overline{z}\mathop{\rm ker}\nolimits T_{g}$, in the context of $L^{2}$, the answer is given by saying that $\mathop{\rm ker}\nolimits T_{g}$ is almost invariant with defect 1 for both $S^{*}$ (in $H^{2}_{+})$ and $M_{\overline{z}}$ (in $L^{2}$). Thus in the latter case one gets information on the image of $\mathop{\rm ker}\nolimits T_{g}$ by $S^{*}$, while in the case of near invariance one gets information describing $\mathop{\rm ker}\nolimits T_{g}$ itself; in particular, there must be an element of $\mathop{\rm ker}\nolimits T_{g}$ which does not vanish at 00, and so $\mathop{\rm ker}\nolimits T_{g}\not\subseteq zH^{2}_{+}$.
In the more general context of $L^{2}$, to extend the previous results to paired operators, one can look at near invariance in the context of $\mathcal{N}$-stability, or stability in $\mathcal{N}$, introduced in [5].
Definition 5.11.
Let $\mathcal{A}\neq\{0\}$ be a closed subspace of $\mathcal{N}\subset\mathcal{H}$ and let $X\in\mathcal{L}(\mathcal{H})$. We say that $\mathcal{A}$ is $\mathcal{N}$-stable for $X$, or stable in $\mathcal{N}$ for $X$, if $X(\mathcal{A})\subset\mathcal{N}$.
Thus, according to this definition, saying that $\mathcal{M}$ is nearly $S^{*}$-invariant is equivalent to saying that $\mathcal{M}\cap zH^{2}_{+}$ is $\mathcal{M}$-stable for the operator $S^{*}$; $\mathcal{M}$ is invariant for $S^{*}$ if and only if it is $\mathcal{M}$-stable for $S^{*}$.
When $\mathcal{M}$ is the kernel of an operator $T$ on $\mathcal{N}\subset\mathcal{H}$, this notion can be related with the commutation relations between two operators as follows. It is clear that if $X$ and $T$ commute then all elements of $\mathop{\rm ker}\nolimits T$ are mapped back into $\mathop{\rm ker}\nolimits T$ by the operator $X$; i.e., $\mathop{\rm ker}\nolimits T$ is invariant for $X$. But if $X$ and $T$ do not commute on their domain, then those elements of $\mathop{\rm ker}\nolimits T$ for which $X$ and $T$ do commute, i.e.,
=$\mathcal{M}=\{f\in\mathop{\rm ker}\nolimits T:XTf=TXf\}$
are also mapped back into $\mathop{\rm ker}\nolimits T$ by $X$, and so $\mathcal{M}$ is $\mathop{\rm ker}\nolimits T$-stable for $X$.
Comparing this with (5.3), note that for $g\not\in\overline{H^{\infty}}$, taking $T=T_{g}$ and $X=S^{*}$, the elements of $\mathcal{M}$, as defined above, are precisely those with $f(0)=0$, so one must indeed have that $S^{*}f\in\mathop{\rm ker}\nolimits T_{g}$. Noting moreover that $\overline{z}f\in H^{2}_{+}$ means that $f\in\mathop{\rm ker}\nolimits H_{\overline{z}}$, we see that in this case
$\mathcal{M}=\mathop{\rm ker}\nolimits T_{g}\cap\mathop{\rm ker}\nolimits H_{\overline{z}}$
and near $S^{*}$-invariance can be expressed as saying that $\mathop{\rm ker}\nolimits T_{g}\cap\mathop{\rm ker}\nolimits H_{\overline{z}}$ is stable in $\mathop{\rm ker}\nolimits T_{g}$ for the operator $S^{*}$. One could express near $\eta$-invariance analogously by saying that $\mathop{\rm ker}\nolimits T_{g}\cap\mathop{\rm ker}\nolimits H_{\eta}$ is stable in $\mathop{\rm ker}\nolimits T_{g}$ for $M_{\eta}$.
Applying these ideas now to the setting of paired operators in $L^{2}(\mathbb{T})$, we have from Propositions 2.12 and 2.13 that, for $\eta\in L^{\infty}$ and $f\in L^{2}$,
$\begin{array}[]{rclcl}\eta S_{a,b}f&=&S_{a,b}\eta f&\iff&f\in\mathop{\rm ker}\nolimits H_{\eta}\oplus\mathop{\rm ker}\nolimits\widetilde{H}_{\eta}\\ \eta\Sigma_{a,b}f&=&\Sigma_{a,b}\eta f&\iff&(a-b)f\in\mathop{\rm ker}\nolimits H_{\eta}\oplus\mathop{\rm ker}\nolimits\widetilde{H}_{\eta}.\end{array}$
Thus we have the following analogue of near $\eta$-invariance.
Proposition 5.12.
(i) $\mathop{\rm ker}\nolimits S_{a,b}\cap(\mathop{\rm ker}\nolimits H_{\eta}\oplus\mathop{\rm ker}\nolimits\widetilde{H}_{\eta})$ is stable in $\mathop{\rm ker}\nolimits S_{a,b}$ for multiplication by $\eta$; i.e.,
$f\in\mathop{\rm ker}\nolimits S_{a,b},\ H_{\eta}f_{+}=0,\ \widetilde{H}_{\eta}f_{-}=0\implies\eta f\in\mathop{\rm ker}\nolimits S_{a,b}.$
(ii) If $f\in\mathop{\rm ker}\nolimits\Sigma_{a,b}$ and $(a-b)f\in\mathop{\rm ker}\nolimits H_{\eta}\oplus\mathop{\rm ker}\nolimits\widetilde{H}_{\eta}$, then $\eta f\in\mathop{\rm ker}\nolimits\Sigma_{a,b}$.
Corollary 5.13.
(i) $f\in\mathop{\rm ker}\nolimits S_{a,b},\ f_{+}\in\mathop{\rm ker}\nolimits H_{\overline{z}}\implies\overline{z}f\in\mathop{\rm ker}\nolimits S_{a,b}$, so there exists in every non-zero paired kernel an element $f$ with $f_{+}(0)\neq 0$.
(ii) $f\in\mathop{\rm ker}\nolimits S_{a,b},\ f_{-}\in\mathop{\rm ker}\nolimits\tilde{H}_{z}\implies zf\in\mathop{\rm ker}\nolimits S_{a,b}$, so there exists in every non-zero paired kernel an element $f$ with $zf_{-}(\infty)\neq 0$.
Note that (i) in the above corollary generalizes a well-known property of nonzero nearly $S^{*}$-invariant subspaces of $H^{2}_{+}$.